OpenStudy (anonymous):

evaluate limit->0 1-secx / 1-cosx?

OpenStudy (anonymous):

can i start with 1-1/cosx / 1-cosx?

OpenStudy (anonymous):

yes that would be a good idea

myininaya (myininaya):

\[\frac{1-\frac{1}{\cos(x)}}{1-\cos(x)} \cdot \frac{\cos(x)}{\cos(x)} =\frac{\cos(x)-1}{\cos(x)(1-\cos(x))}\]

myininaya (myininaya):

what do you think is next

OpenStudy (anonymous):

but then i dont know what to do next

myininaya (myininaya):

cos(x)-1 is the opposite of 1-cos(x) right?

myininaya (myininaya):

\[\frac{\cos(x)-1}{\cos(x)(-1)(\cos(x)-1)}\]

OpenStudy (anonymous):

yes?

myininaya (myininaya):

so recall a/a=1 assuming a does not =0

OpenStudy (anonymous):

nice okay i got it = -1!!

myininaya (myininaya):

ok great

OpenStudy (anonymous):

yes that it is!

OpenStudy (anonymous):

ok im probably going to post again soon if i get stuck with more of these hehe..